ADNAN MUHAMMED

Grade 12,

(a) Statement-1 is True, Statement-2 is True; Statement-2 is a correct explanation for Statement-1. (b) Statement-1 is True, Statement-2 is True; Statement-2 is NOT a correct explanation for Statement-1. (c) Statement-1 is True, Statement-2 is False. (d) Statement-1 is False, Statement-2 is True. Statement-1: For projection angle tan –1 (4), the horizontal range and the maximum height of a projectile are equal. Statement-2: The maximum range of projectile is directly proportional to square of velocity and inversely proportional to acceleration due to gravity.



(a)    Statement-1 is True, Statement-2 is True; Statement-2 is a correct explanation for Statement-1.


(b)    Statement-1 is True, Statement-2 is True; Statement-2 is NOT a correct explanation for Statement-1.


(c)    Statement-1 is True, Statement-2 is False.


(d)    Statement-1 is False, Statement-2 is True.


 






Statement-1: For projection angle tan–1(4), the horizontal range and the maximum height of a projectile are equal.


       


 Statement-2: The maximum range of projectile is directly proportional to square of velocity and inversely proportional to acceleration due to gravity.



 

Grade:upto college level

1 Answers

Simran Bhatia
10 years ago

(b)

    The horizontal range of projectile, R = (u^2  sin?2x)/g

    Rmax = u^2/g

    The maximum height attained by projectile

    H = (u^2 sin^2 x)/2g

    If H = R (u^2 sin^2 x)/2g=(u^2 sin 2x)/g

    (sin^2 x)/2 = 2 sinx cosx =>tanx = 4

    x= tan–1 (4)